The locus of the point of intersection of the lines y+mx=(a2m2+b2) and my-x=√ (a2+b2m2)is??[m is a parameter]

Grade:11

1 Answers

Dilli PRASAD
6 years ago
Square the given eqts Y^2+m^2x^2+2ymx= a^2m^2+b^2m^2y^2+x^2-2mxy=a^2+b^2m^2By adding above eqts Y^2(1+m^2)+x^2(1+m^2)=(a^2+b^2)(1+m^2)There fore locus is x^2+y^2=a^2+b^2